The current price of YBM stock S is $101. American options with a strike price K = $100 and maturing in T = 6 months trade on YBM. The continuously compounded, risk-free interest rate r is 5 percent per year. If the American put price pA is $2.70, then the American call price cA will at maximum be:
Answer #1

I have answered the question below w.

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

According to the Put-call parity;

C0 + X*e-r*t = P0 + S0

C0 + [$100 * e-[0.05 * (6/12)]] = $2.70 + $101

C0 + [$100 * 0.9753] = $103.70

C0 + $97.53 = $103.70

C0 = $103.70 - $97.53 = $6.17